# Probing Geometrical NSI at the DUNE experiment

### v1

@@ Abstract
In this work, we investigate the implications of a novel non-standard interaction (NSI) of neutrinos. This interaction is geometric in origin – it arises because the propagation of fermions in curved spacetime induces torsion. This torsion is non-propagating and can be eliminated from the action, resulting in a four-fermion interaction in a torsion-free background. The new interaction modifies the behaviour of the neutrinos passing through matter by introducing additional coupling terms, resulting in a new component in the effective potential. As a result, the neutrino oscillation probabilities in matter are altered. The relevant probabilities are computed using the Cayley-Hamilton formalism. We then explore the potential to probe these torsion-induced NSI in the proposed DUNE experiment. Constraints on the parameters characterizing the torsional effects are obtained. By selecting representative values of torsion parameters to which the DUNE experiment is sensitive, we analyse how these geometric interactions affect the experiment’s sensitivity to determine neutrino mass hierarchy, the octant of the 2-3 leptonic mixing angle, and the CP phase. We also examine the new parameter degeneracies introduced by torsion effects and assess their impact on the overall sensitivities of DUNE. We find that the additional parameter degeneracies in the presence of torsion significantly affect the octant sensitivity.
@@ I Introduction
The discovery of neutrino oscillations, where neutrinos transform between different flavour states as they propagate, has unequivocally established that neutrinos have mass. Since neutrinos are massless in the Standard Model of particle physics, the observation of neutrino oscillation in several experiments also provided the first compelling evidence of physics beyond the Standard Model (BSM). The past and ongoing experiments [1, 2, 3, 4, 5, 6, 7] have also determined a majority of the parameters governing the oscillation of neutrinos with considerable precision. These are the mass squared differences $\Delta m^{2}_{21}$ and $|\Delta m^{2}_{31}|$, where $\Delta m^{2}_{ij}=m_{i}^{2}-m_{j}^{2}$ , and the mixing angles $\theta_{12}$ , $\theta_{23}$ , and $\theta_{13}$. Currently, the unknown parameters are the neutrino mass ordering or the sign of $\Delta m^{2}_{31}$, the octant of $\theta_{23}$, and the value of the leptonic CP phase, $\delta_{CP}$. The mass ordering or hierarchy can be normal (NH), in which $m_{1}<m_{2}<m_{3}$ or inverted (IH) when $m_{3}<m_{1}<m_{2}\,$, whereas the angle $\theta_{23}$ can be $45^{o}$ i.e in the lower octant (LO) or $45^{o}$ i.e in the higher octant (HO). One of the main reasons these parameters remain undetermined by current generation experiments is the existence of degeneracies, which imply different combinations of oscillation parameters that yield the same value of the probability [8, 9, 10, 11]. Specifically, it was shown in Ref. [12] that the generalized hierarchy-$\theta_{23}$-$\delta_{CP}$ degeneracy impairs the capability of the currently running experiments in precisely determining the above three oscillation parameters. Many new high-intensity and high-statistics experiments are underway, and these are expected to resolve the degeneracies leading to an unambiguous determination of the hierarchy, octant of $\theta_{23}$, and $\delta_{CP}$. These include the beam-based experiments DUNE [13], T2HK [14], reactor-based experiments JUNO [15] etc.
@@
In this work, we examine the implications of a special type of non-standard interaction, which arises when fermions propagate through a background of fermionic matter and is mediated by torsion in curved spacetime [16]. This interaction modifies the dynamics of the fermions, introducing additional coupling terms, giving rise to an effective potential that depends on the density and spin of the background matter. A derivation of the potential at finite temperature and densities is given in Ref. [17]. In the case of neutrinos, this effect becomes especially interesting, as it can lead to observable deviations in their oscillation behaviour.
We investigate how the presence of these new interactions originating from the geometry of spacetime can impact the neutrino propagation through a distance of $\sim 1300$ km, which is the proposed baseline for the DUNE experiment. We determine the constraints on the parameters governing this effect from an analysis of simulated DUNE data. In addition, we probe how the presence of these new interactions affects the capability of the DUNE experiment to determine the hierarchy, octant, and CP phase. We consider the interplay of the disappearance and appearance channels in determining these parameters and also discuss how the occurrence of degeneracies affects the determination of the parameters in the presence of this new effect.
@@
The effect of spacetime torsion on neutrino oscillations has been investigated before [18, 19, 20, 21]. However, what distinguishes the scenario considered in this work is that here the torsion is not a field in the background spacetime, but is dynamically generated by the fermions themselves. The torsion acts as an auxiliary field and gives rise to an effective four-fermion interaction, which can be considered as chiral and non-universal. The computation of the neutrino oscillation probabilities in the presence of such fermion-induced torsion has been performed for matter of constant density [22], in the context of atmospheric neutrinos [23] as well as for Supernova neutrinos [24]. An earlier study in the context of DUNE and P2SO experiments can be found in Refs. [22, 25].
@@
The plan of the paper is as follows. In Sec. II we outline the analytic formulation of computing the probabilities in the presence of torsion. In Sec. III we present plots of the oscillation probabilities which bring out the effect of torsion, and discuss how torsion affects the hierarchy, octant, and CP sensitivities. Sec. IV presents the analysis methodology and the details of the experimental setup used in our analysis, as well as the results. We end with a summary of our results.
@@ II Neutrino oscillations in presence of torsion
$H_{I}=\sum_{i=e,\mu,\tau}\left(\lambda_{i}{\nu}_{iL}^{\dagger}\nu_{iL}\right)\,\tilde{n}\,$, (1)
where $\tilde{n}$ is the weighted density of the background fermions,
@@
The coupling constants $\lambda\,$, which have dimensions of length, cannot be determined from theoretical considerations, but must be fixed by experiments. We write $\lambda_{i}$ for the couplings corresponding to $\nu_{i}$ , and $\lambda^{f}_{V,A}$ for those corresponding to the vector (axial) current of the fermion $f$ . We will refer to the $\lambda$ as torsional or geometrical coupling constants.
@@
where $c_{ij}=\cos\theta_{ij}$ and $s_{ij}=\sin\theta_{ij}\,.$ The angles $\theta_{ij}$ can be taken to lie in the first quadrant, while $\delta$ varies between $180^{\circ}$ and $180^{\circ}$  while $\eta_{1},\eta_{2}$ can take values in the range between 0 and $180^{\circ}$. We will consider Dirac neutrinos, so that the Majorana phases can be absorbed by a redefinition of fields. Then the mixing matrix can be conveniently expressed as a product of rotation matrices ${\mathcal{O}}_{ij}$ for rotation in the $ij$-plane and $U_{\delta}=\mathrm{diag}(1,\,1,\,e^{i\delta})$ ,
@@
$i\frac{d}{dt}{\begin{pmatrix}{\nu_{e}}\\ {\nu_{\mu}}\\ {\nu_{\tau}}\end{pmatrix}}$ $=\frac{1}{2E}\left[U\left[\begin{pmatrix}0~&0&0\\ 0~&\Delta{m}_{12}^{2}&0\\ 0~&0&\Delta{m}_{31}^{2}\end{pmatrix}+2E\tilde{n}\begin{pmatrix}0&0&0\\ 0&\lambda_{21}&0\\ 0&0&\lambda_{31}\end{pmatrix}\right]U^{\dagger}+\begin{pmatrix}A\Delta m^{2}_{31}&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}\right]\,{\begin{pmatrix}{\nu_{e}}\\ {\nu_{\mu}}\\ {\nu_{\tau}}\end{pmatrix}}\,$. (6)
@@
Based on global fits to neutrino data [34], we know that $\alpha\approx\pm 0.03$, while $s_{13}\approx 0.15$. Also, if we assume $\lambda_{e}=\lambda_{u}=\lambda_{d}=\lambda_{ij}=0.1\sqrt{G_{F}}$, we get $\beta_{ij}\approx 0.0075$ at $E=2.5$ GeV. In our perturbative calculation, we assume $\alpha,~s_{13},~\beta_{21},~\beta_{31}$ to be small parameters, and we keep terms up to the second order in them. Note that the value of $\beta_{ij}$ depends on energy. For $E\approx 10$ GeV, we get $\beta_{ij}\sim\alpha$ (see Appendix B for details), while for lower energies $\beta_{ij}<\alpha$ .
@@
The corresponding expressions for antineutrinos can be obtained by replacing $\delta_{CP}\to-\delta_{CP}$, $A\to-A$, and $\beta_{ij}\to-\beta_{ij}$.
@@ IV.1 Experimental setup and numerical analysis details
We consider a liquid Argon Detector similar to what is proposed by the DUNE collaboration -power The GLoBES package [36, 37] has been used to simulate the DUNE setup. The values of interaction cross-sections, energy resolutions, efficiencies, backgrounds, and systematic errors are taken from the DUNE collaboration [38]. These include Geant-simulated neutrino fluxes, cross-sections included from GENIE, and smearing matrices and post-smearing efficiencies to calculate the expected event rates.
@@ IV.2 Bounds on torsion parameters
Figure 4: Projected bounds on $\lambda_{(2,3)1}$ from DUNE. The (left) right panel shows the bounds on the geometrical couplings with the (normal) inverted hierarchy. The top panels show the bounds without any marginalization in the test spectrum. The bottom panels show the bounds with marginalization over $\delta_{CP},\theta_{23},\Delta m_{31}^{2}$ in the test spectrum.
In Fig. 4 the constraints on the torsion parameters $\lambda_{21}$ and $\lambda_{31}$ are presented considering disappearance and appearance channels separately and also combining both. The plots in the top row represent the bounds for fixed parameters while for the plots in the bottom row, the effect of parameter uncertainties are included through marginalization over $\theta_{23}$, $\delta_{CP}$ and $|\Delta m^{2}_{31}|$. The sign of $\Delta m^{2}_{31}$ is kept fixed. It is seen that the appearance channel constrains $\lambda_{21}$ better than $\lambda_{31}$. This can be understood from Eq. (10) which shows that the appearance probability depends on $\beta_{21}$ but not on $\beta_{31}$ up to the second order. (We remind the reader that the $\beta_{ij}$ parameters scale linearly with $\lambda_{ij}$ for given background torsional couplings, as given in Eq. (8).)
@@ IV.3 Hierarchy
  • The mass hierarchy sensitivity is very high for all values of $\delta_{CP}$ for all the cases. The lowest sensitivity occurs at $\delta_{CP}=+90^{o}(-90^{o})$ for NH (IH) and is $\sim 10\sigma$.

### v2

@@ Abstract
In this work, we investigate the implications of a novel non-standard interaction (NSI) of neutrinos. This interaction is geometric in origin – it arises because the propagation of fermions in curved spacetime induces torsion. This torsion is non-propagating and can be eliminated from the action, resulting in a four-fermion interaction in a torsion-free background. The new interaction modifies the behaviour of the neutrinos passing through matter by introducing additional coupling terms, resulting in a new component in the effective potential. As a result, the neutrino oscillation probabilities in matter are altered. The relevant probabilities are computed using the Cayley-Hamilton formalism. We then numerically explore the potential to probe these torsion-induced NSI in the DUNE experiment. We obtain the bounds on the parameters characterizing the torsional effects. By selecting representative values of torsion parameters to which the DUNE experiment is sensitive, we analyse how these geometric interactions affect the experiment’s sensitivity to determine neutrino mass hierarchy, the octant of the 2-3 leptonic mixing angle, and the CP phase. We also examine the new parameter degeneracies introduced by torsion effects and assess their impact on the overall sensitivities of DUNE. We find that the additional parameter degeneracies in the presence of torsion significantly affect the octant sensitivity.
@@ I Introduction
The discovery of neutrino oscillations, where neutrinos transform between different flavour states as they propagate, has unequivocally established that neutrinos have mass. Since neutrinos are massless in the Standard Model of particle physics, the observation of neutrino oscillation in several experiments also provided the first compelling evidence of physics beyond the Standard Model (BSM). The past and ongoing experiments [1, 2, 3, 4, 5, 6, 7] have also determined a majority of the parameters governing the oscillation of neutrinos with considerable precision. These are the mass squared differences $\Delta m^{2}_{21}$ and $|\Delta m^{2}_{31}|$, where $\Delta m^{2}_{ij}=m_{i}^{2}-m_{j}^{2}$ , and the mixing angles $\theta_{12}$ , $\theta_{23}$ , and $\theta_{13}$. Currently, the unknown parameters are the neutrino mass ordering or the sign of $\Delta m^{2}_{31}$, the octant of $\theta_{23}$, and the value of the leptonic CP phase, $\delta_{CP}$. The mass ordering or hierarchy can be normal (NH), in which $m_{1}<m_{2}<m_{3}$ or inverted (IH) when $m_{3}<m_{1}<m_{2}\,$, whereas the angle $\theta_{23}$ can be $45^{\circ}$ i.e in the lower octant (LO) or $45^{\circ}$ i.e in the higher octant (HO). One of the main reasons these parameters remain undetermined by current generation experiments is the existence of degeneracies, which imply different combinations of oscillation parameters that yield the same value of the probability [8, 9, 10, 11]. Specifically, it was shown in Ref. [12] that the generalized hierarchy-$\theta_{23}$-$\delta_{CP}$ degeneracy impairs the capability of the currently running experiments in precisely determining the above three oscillation parameters. Many new high-intensity and high-statistics experiments are underway, and these are expected to resolve the degeneracies leading to an unambiguous determination of the hierarchy, octant of $\theta_{23}$, and $\delta_{CP}$. These include the beam-based experiments DUNE [13], T2HK [14], reactor-based experiments JUNO [15] etc.
@@
In this work, we examine the implications of a special type of non-standard interaction, which arises when fermions propagate through a background of fermionic matter and is mediated by torsion in curved spacetime [16]. This interaction modifies the dynamics of the fermions, introducing additional coupling terms, giving rise to an effective potential that depends on the density and spin of the background matter. A derivation of the potential at finite temperature and densities is given in Ref. [17]. In the case of neutrinos, this effect becomes especially interesting, as it can lead to observable deviations in their oscillation behaviour. We investigate how the presence of these new interactions originating from the geometry of spacetime can impact the neutrino propagation through a distance of $\sim 1300$ km, which is the proposed baseline for the DUNE experiment. We determine the constraints on the parameters governing this effect from an analysis of simulated DUNE data. In addition, we probe how the presence of these new interactions affects the capability of the DUNE experiment to determine the hierarchy, octant, and CP phase. We consider the interplay of the disappearance and appearance channels in determining these parameters and also discuss how the occurrence of degeneracies affects the determination of the parameters in the presence of this new effect.
Different kinds of theories are possible on a Riemann-Cartan spacetime, which is the broad name for spacetimes with torsion [18]. In Poincaré gauge theory, translations and Lorentz transformations are both gauged. Tetrads and the spin connection are the corresponding gauge fields; in general the theory contains propagating torsion as part of the spin connection. Teleparallel gravity describes theories in which the connection is flat but contains torsion [19, 20, 21, 22]; torsion is a propagating field in these theories as well. Theories in which torsion is sourced by fermions are called Einstein-Cartan theories; we further use the Sciama-Kibble formulation (ECSK theory) [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] and allow a chiral coupling of fermions to torsion.
@@
The effect of spacetime torsion on neutrino oscillations has been investigated before [34, 35, 36, 37, 38, 39]. However, what distinguishes the scenario considered in this work is that here the torsion is an auxiliary field, which couples chirally and non-universally to fermions. In other words, the torsion is not a propagating field in the background spacetime, but is dynamically generated by the fermions themselves. The torsion acts as an auxiliary field and gives rise to an effective four-fermion interaction, which can be considered as chiral and non-universal – see Appendix A. The computation of the neutrino oscillation probabilities in the presence of such fermion-induced torsion has been performed for matter of constant density [40], in the context of atmospheric neutrinos [41] as well as for Supernova neutrinos [42]. An earlier study in the context of DUNE and P2SO experiments can be found in Refs. [40, 43].
@@
The plan of the paper is as follows. In the next section, we outline the analytic formulation of computing the probabilities in presence of torsion. In Sec. III we present plots of the oscillation probabilities which bring out the effect of torsion, and discuss how torsion affects the hierarchy, octant, and CP sensitivities. Sec. IV presents the analysis methodology and the details of the experimental setup used in our analysis, as well as the results. We end with a summary of our results.
@@ II Neutrino oscillations in presence of torsion
$H_{I}=\sum_{i=1,2,3}\left(\lambda_{i}{\nu}_{iL}^{\dagger}\nu_{iL}\right)\,\tilde{n}\,$, (1)
where the sum runs over fermions in the mass basis. Since this interaction arises purely from spacetime, independently of any other interaction, it is naturally diagonal in the mass basis. Although it is possible in principle for the torsional interaction to be diagonal in a different ‘torsion’ basis, such a choice will introduce additional mixing matrices with several unknown parameters. We will not consider that possibility in this paper. Also, here $\tilde{n}$ is the weighted density of the background fermions,
@@
The coupling constants $\lambda\,$, which have dimensions of length, cannot be determined from theoretical considerations, but must be fixed by experiments. This is because the ECSK theory is based on Poincaré gauge theory, with a gauge group which is the semi-direct product of non-compact Lorentz and translation groups, which will have independent coupling constants. Thus gravity and torsion couple to fields with different couplings and there is no relation between those coupling constants. In particular, while the scale of gravitational interaction is determined by its classical limit, there is no classical limit for the spin-torsion interaction, so its scale cannot be determined by that method. We write $\lambda_{i}$ for the couplings corresponding to $\nu_{i}$ , and $\lambda^{f}_{V,A}$ for those corresponding to the vector (axial) current of the fermion $f$ . We will refer to the $\lambda$ as torsional or geometrical coupling constants. The background fermions should also be in the mass basis. However, we will assume that all the background fermions have approximately the same values of the torsional coupling $\lambda\,$, and thus will ignore the flavor mixing of $d$-type quarks for the purpose of this paper.
@@
where $c_{ij}=\cos\theta_{ij}$ and $s_{ij}=\sin\theta_{ij}\,.$ The angles $\theta_{ij}$ can be taken to lie in the first quadrant, while $\delta$ varies between $180^{\circ}$ and $180^{\circ}$  and $\eta_{1},\eta_{2}$ can take values in the range between 0 and $180^{\circ}$. We will consider Dirac neutrinos, so that the Majorana phases can be absorbed by a redefinition of fields. Then the mixing matrix can be conveniently expressed as a product of rotation matrices ${\mathcal{O}}_{ij}$ for rotation in the $ij$-plane and $U_{\delta}=\mathrm{diag}(1,\,1,\,e^{i\delta})$ ,
@@
$i\frac{d}{dt}{\begin{pmatrix}{\nu_{e}}\\ {\nu_{\mu}}\\ {\nu_{\tau}}\end{pmatrix}}$ $=\frac{1}{2E}\left[U\left[\begin{pmatrix}0~&0&0\\ 0~&\Delta{m}_{12}^{2}&0\\ 0~&0&\Delta{m}_{31}^{2}\end{pmatrix}+2E\tilde{n}\begin{pmatrix}0&0&0\\ 0&\lambda_{21}&0\\ 0&0&\lambda_{31}\end{pmatrix}\right]U^{\dagger}+\begin{pmatrix}A\Delta m^{2}_{31}&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}\right]\,{\begin{pmatrix}{\nu_{e}}\\ {\nu_{\mu}}\\ {\nu_{\tau}}\end{pmatrix}}\,$, (6)
@@
Based on global fits to neutrino data [52], we know that $\alpha\approx\pm 0.03$, while $s_{13}\approx 0.15$. Also, if we assume $\lambda_{e}=\lambda_{u}=\lambda_{d}=\lambda_{ij}=0.1\sqrt{G_{F}}$ (see Sec. IV.1), we get $\beta_{ij}\approx 0.0075$ at $E=2.5$ GeV. In our perturbative calculation, we assume $\alpha,~s_{13},~\beta_{21},~\beta_{31}$ to be small parameters, and we keep terms up to the second order in them. Note that the value of $\beta_{ij}$ depends on energy. For $E\approx 10$ GeV, we get $\beta_{ij}\sim\alpha$ (see Appendix B for details), while for lower energies $\beta_{ij}<\alpha$ .
@@
The corresponding expressions for antineutrinos can be obtained by replacing $\delta_{CP}\to-\delta_{CP}$, $A\to-A$, and $\beta_{ij}\to~-\beta_{ij}$.
The geometrical interaction is analogous to the neutral-current-mediated interaction known commonly in the literature as NSI, but there are some key differences. The geometrical interaction is diagonal in the mass basis, while the ‘usual’ NSIs are expressed in the flavour basis. It is of course simple to map one set of new physics parameters to the other. However, in this case the elements of the geometrical Hamiltonian in the flavour basis will not all be independent, since the relevant physics is described by only two parameters – $\lambda_{21}$ (or $\beta_{21}$) and $\lambda_{31}$ (or $\beta_{31}$). Thus the key difference between ‘usual’ NSIs and geometrical NSIs in flavour basis is the existence of correlations between the various elements.
Another difference, which is very important from a physics perspective, is that the geometrical interaction term is proportional to a weighted matter density. This is because the geometrical interaction between any two fermions is proportional to the product of the torsional coupling constants associated with the two fermions. In the usual NC-NSI models, the interaction terms between the background and the neutrinos in the flavour basis are included with arbitrary parameters. It is of course possible to choose the parameters to be related in such a way that they describe the geometrical NSI, but arbitrary parameters do not originate from a physical model. Thus, the geometrical NSI can be seen as a special case of all possible NC-NSI.
@@ IV.1 Experimental setup and numerical analysis details
We consider a 40 kiloton liquid argon detector similar to what is proposed by the DUNE collaboration. The experimental setup is simulated using the GLoBES package [54, 55]. In doing so, the neutrino fluxes, interaction cross-sections, energy resolutions and smearing matrices, selection efficiencies, backgrounds, and systematic errors are taken from the DUNE collaboration [56]. The event rates generated using the above inputs are used for our statistical analyses.
@@
We have taken the $\lambda$ values of the background fermions to be (of the order) $0.1\sqrt{G_{F}}$ . This number is within the bound estimated in [57] from parity violation in electron-electron and electron-deuteron scattering. Furthermore, we have assumed $\lambda_{e}=\lambda_{u}=\lambda_{d}\,$ in order to simplify calculations.
@@ IV.2 Bounds on torsion parameters
Figure 4: Projected $3\sigma$ bounds on $\lambda_{(2,3)1}$ from DUNE. The (left) right panel shows the bounds on the geometrical couplings with the (normal) inverted hierarchy. The top panels show the bounds without any marginalization in the test spectrum. The bottom panels show the bounds with marginalization over $\delta_{CP},\theta_{23},\Delta m_{31}^{2}$ in the test spectrum.
In Fig. 4 the $3\sigma$ constraints on the torsion parameters $\lambda_{21}$ and $\lambda_{31}$ are presented considering disappearance and appearance channels separately and also combining both. The plots in the top row represent the bounds for fixed parameters while for the plots in the bottom row, the effect of parameter uncertainties are included through marginalization over $\theta_{23}$, $\delta_{CP}$ and $|\Delta m^{2}_{31}|$. The sign of $\Delta m^{2}_{31}$ is kept fixed. It is seen that the appearance channel constrains $\lambda_{21}$ better than $\lambda_{31}$. This can be understood from Eq. (10) which shows that the appearance probability depends on $\beta_{21}$ but not on $\beta_{31}$ up to the second order. (We remind the reader that the $\beta_{ij}$ parameters scale linearly with $\lambda_{ij}$ for given background torsional couplings, as given in Eq. (8).)
@@ IV.3 Hierarchy
  • The mass hierarchy sensitivity is very high for all values of $\delta_{CP}$ for all the cases. The lowest sensitivity occurs at $\delta_{CP}=+90^{\circ}(-90^{\circ})$ for NH (IH) and is $\sim 10\sigma$.
@@
Table 2 summarizes the degeneracies between the four hierarchy and octant combinations in the parameter space, for neutrinos and antineutrinos in the presence of positive, zero, and negative torsion. Similarly, Table 3 summarizes the degeneracies between the four hierarchy and $\delta_{CP}$ half-plane combinations. Degeneracies that exist only for neutrinos or only for antineutrinos can be easily lifted by collecting data with both polarities.
Parameters NH LO NH HO IH LO IH HO
No torsion $\nu$ Degenerate with IH HO, IH LO No degeneracy Degenerate with NH LO Degenerate with NH LO
Negative torsion $\nu$ Less degeneracy No degeneracy No degeneracy Less degeneracy
Positive torsion $\nu$ More degeneracy Degenerate with IH HO More degeneracy More degeneracy
No torsion $\bar{\nu}$ No degeneracy Degenerate with IH LO Degenerate with NH HO No degeneracy
Negative torsion $\bar{\nu}$ Degenerate with IH HO More degeneracy More degenerate with NH HO, Degenerate with NH LO Degenerate with NH HO
Positive torsion $\bar{\nu}$ No degeneracy No degeneracy No degeneracy No degeneracy
Table 2: The hierarchy-octant degenerate parameter space for neutrinos and antineutrinos. Here, LO = Lower octant, HO = Higher octant, NH = Normal hierarchy, IH = Inverted hierarchy. The positive(negative) torsion takes $\lambda_{(2,3)1}=\pm 0.1\sqrt{G_{F}}$.
Parameters NH LHP NH UHP IH LHP IH UHP
No torsion $\nu$ No degeneracy Degenerate with IH LHP Degenerate with NH UHP No degeneracy
Negative torsion $\nu$ No degeneracy No degeneracy No degeneracy No degeneracy
Positive torsion $\nu$ No degeneracy More degeneracy More degeneracy No degeneracy
No torsion $\bar{\nu}$ No degeneracy Degenerate with IH LHP Degenerate with NH UHP No degeneracy
Negative torsion $\bar{\nu}$ No degeneracy More degeneracy More degeneracy No degeneracy
Positive torsion $\bar{\nu}$ No degeneracy No degeneracy No degeneracy No degeneracy
Table 3: The hierarchy-$\delta_{CP}$ degenerate parameter space for neutrinos and antineutrinos. Here, LO = Lower octant, HO = Higher octant, NH = Normal hierarchy, IH = Inverted hierarchy. The positive(negative) torsion takes $\lambda_{(2,3)1}=\pm 0.1\sqrt{G_{F}}$.